Circular No. 7810

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ap IN M74 S. Nakano, Sumoto, Japan, reports the discovery by Yoji Hirose, Chi-

gasaki, Kanagawa-ken, of an apparent supernova (V = 14.5) on five unfiltered CCD frames (limiting mag 17) taken on Jan. 29.398–29.501 UT with a 0.25-m f/6.3 Schmidt-Cassegrain reflector. Hirose reported the new object as being 274′′ west and 93′′ south of the center of M74 (= NGC 628). Nothing was visible at this location on his CCD frames taken on Jan. 3. R. Kushida, Yatsugatake South Base Observatory, reports the following precise position (measured by Y. Kushida) for SN 2002ap from an unfiltered CCD image taken on Jan. 30.382, with the star at mag 13.7: α = 1^{h}36^{m}23^{s}.85, δ = +15^{o}45′13′′.2 (equinox 2000.0), or 258′′ west and 108′′ south of the center of M74. No star is visible at this position on Kushida’s survey frames taken from 1994 to 2001 Dec. 18 (limiting mag 17–18). W. Li, University of California, reports that an image taken with the Katzman Automatic Imaging Telescope (KAIT) on Jan. 29.2 showed the new object at mag 14.4. Li also measured position end figures for SN 2002ap: 23^{s}.87, 13′′.0. A KAIT image taken on Jan. 25.2 showed nothing at this position (limiting mag ∼ 18.0).

V2540 OPHIUCHI = NOVA OPHIUCHI 2002 N. N. Samus, Institute of Astronomy, Moscow, informs us that the

designation V2540 Oph has been given to this nova (cf. IAUC 7808, 7809).

SUPERNOVA 2002ao IN UGC 9299 A. Gal-Yam and O. Shemmer, Tel Aviv University, write: “A spectrum

(range 400–780 nm) of SN 2002ao (IAUC 7809), obtained under poor conditions with the Wise Observatory 1-m telescope (+ FOSC spectrograph) on Jan. 30 by J. Dann, shows a blue continuum with broad features. The Si II trough typical of type-Ia supernovae is not present, neither is the strong He I line around 570 nm that is observed in type-Ib supernovae. The object is bluer than any of the spectra of type-Ic events that are available to us (e.g., Filippenko 1997, Ann. Rev. A.Ap. 35, 309), so we think it unlikely to be of that type. The lack of a strong Hα emission component (either broad or narrow) seems not to favor possible identification of this supernova as of type II-P or IIn. However, we find that the observed spectrum is similar to a spectrum of the type-IIb supernova 1996cb (Qiu et al. 1999, A.J. 117,

736) taken two days prior to maximum light. It therefore seems most likely that SN 2002ao is of type IIb, observed close to peak brightness.”

2002 January 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green